# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 3.1.5. Summary of Key Information Extraction
Several models like Chargrid (Katti et al., 2018) and ACP (Palm et al., 2019) have enhanced VRDU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilized by (Zhang et al., 2020; Wang et al., 2021a), aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialized models, which produce less representative features compared to those generated by large-scale pre-trained models like LayoutLM (Xu et al., 2020a) and SelfDoc (Li et al., 2021b). Documents typically exhibit specific layouts and logical structures, which has prompted many models (Yu et al., 2021; Lee et al., 2022b, 2023) to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot (Wang and Shang, 2022) and zero-shot (Wang et al., 2023a) frameworks, which leverage contrastive learning and innovative attention mechanisms. Additionally, prompt learning has been applied to distill implicit knowledge from large-scale layout-aware pre-trained models (Hong et al., 2021; Tu et al., 2023) and large language models (LLMs/MLLMs) (He et al., 2023; Perot et al., 2023). Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRDU optimization.
3.2. Document Entity Linking
@@
3.2.1. Entity-Level Entity Linking
@@
DocStruct (Wang et al., 2020b) is the first entity linking framework by leveraging multimodal features of known document entities to predict the hierarchical structure between them. [CLS] token from a BERT-like pretrained language model is extracted for entity-level textual representation $T_{e}$, and an RNN is applied to encode the sequential RoI visual cues $V_{e}$ along the width of the ResNet-50 feature maps. Then, $T_{e}$ is concatenated with layout feature $P_{e}$, which is a linear projected vector of entity bounding box coordinates, $[x1,y1,x2,y2,x3,y3,x4,y4]$. The final entity representation $E$ can be represented:
(4) $\alpha=Sigmoid(W[T_{e},P_{e},V_{e}]+b),$
(5) $E=[T_{e};P_{e}]+\alpha V_{e},$
@@
where $\alpha$ is used as a gate to control the influence of visual cues $V_{e}$ following (Wang et al., 2019). Then, for pair of entities $E_{i}$ and $E_{j}$, the probability of an existing parent-child relationship is represented as $P_{i\rightarrow j}=E_{i}ME_{j}$ where M is the matrix of asymmetric parameters to ensure the asymmetric relation between $E_{i}$ and $E_{j}$. Negative sampling (Mikolov et al., 2013b) is applied during training to handle data sparsity and balance problems.
@@
Zhang et al. (Zhang et al., 2021) propose an entity-linking framework, SERA (Semantic Entity Relation extraction As dependency parsing) by formulating the entity relation prediction as a dependency parsing problem. LayoutLM (Xu et al., 2020a) is used to obtain textual representations at the entity level $T_{e}$ that are concatenated with the embedding of the linear projected entity label $L_{e}$ to obtain the final entity representation $E$, which could be formulated as $E=[T_{e},L_{e}]$. Various encoders, such as Vanilla Transformer, BiLSTM, and Graph, are adopted to learn the following contextual representations. Then, a Biaffine parser is adopted to calculate the score between $E_{i}$ and $E_{j}$. The biaffine parser score $p_{b}$ can be calculated by:
(6) $h_{i}^{key}=\sigma(W^{key}E_{i}+b^{key}),$
(7) $h_{j}^{value}=\sigma(W^{value}E_{j}+b^{value}),$
(8) $p_{b}=h_{i}^{key}W_{b1}h_{j}^{value}+h_{i}^{key}W_{b2}.$
where $h_{i}^{key}$ and $h_{j}^{value}$ are the hidden states after linear projection. To leverage the layout information, a two-dim layout vector $l_{ij}$ is used to get the layout feature score $p_{l}=W_{l}l_{ij}+b_{l}$, where $l_{ij}$ is the minimum distance between two entities along with width and height direction respectively. Then, the final score $p=p_{b}+p_{l}$ is used to calculate loss using binary or multi-label classification.
KVPFormer (Hu et al., 2023) formulate the entity linking to a question-answer problem by introducing a transformer-based encoder-decoder structure to leverage joint-grained information (token and entity) to predict the entity association. The textual representation $T$ of each document entity, $E$, is encoded by averaging the inner-token embedding from pretrained document understanding models, concatenating with the linear projected entity label embedding $l$, represented as $E=[T;l]$. Then, the entity representations will be fed into a Transformer encoder, which has a spatial compatibility attention bias into vanilla self-attention mechanism  [^2] Please refer to . A binary classifier determines which entity is a key entity or not. All detected key entities will be treated as question representation and fed into a DETR-style (Carion et al., 2020) decoder to predict corresponding answers in parallel. For each input question (key query), Top K answer candidates will be selected based on the $Sigmoid$ score and fed into a softmax layer to get the final prediction, named a coarse-to-fine answer prediction.
3.2.2. Token-level Entity Linking
@@
As acquiring entity information needs prior knowledge from manual annotation or layout analysing models, some works utilise serialised OCR-extracted tokens as inputs to extract the structured relations. However, as the logical relation links semantic entities, token-level frameworks must group tokens into entities before exploring their association.
@@
Carbonell et al. (Carbonell et al., 2021) introduces a framework comprising three modules for sequentially conducting token grouping, entity labelling and relation prediction. Firstly, each text token $t$ is represented by $[L_{t};T_{t}]$ where $L_{t}=[x,y,w,h]$ is the coco format bounding box coordinates of $t$ and $T_{t}$ is the work/representations. All tokens are fed into a token grouping GNN, $\mathcal{G}_{group}$, as a node where the edge between nodes is determined by k-NN to avoid high consumption of fully connected GNN. The $\mathcal{G}_{group}$ is trained on a link prediction task to predict the edge score between two nodes to group words, of which scores larger than a predefined threshold $\rho$. Then, the grouped words are fed into a Graph Attention Network (GAT) to use multi-head attention to aggregate words into entities and follow an MLP to conduct node classification to predict the category of each document entity. At last, another link prediction GNN, $\mathcal{G}_{link}$, is trained on edge classification based on aggregated entities.
@@
SPADE (Hwang et al., 2021) formulate the token-level entity linking as a spatial dependency parsing task for serializing (ordering and grouping) tokens and predicting inter-group relation between grouped tokens. Firstly, a spatial text encoder is designed to make spatial-aware attention by introducing a relative spatial vector considering relative, physical and angle aspects. During this task, two binary matrices must be predicted $M_{g}$ for token grouping and inter-group linking $M_{l}$. The vertices comprised by a entity types $\mathbb{V}$ and sequence of tokens $\mathbb{T}$ the encoded entity category and token are represented as $c$ and $t$, respectively, and the relation score between vertices $v_{i}\rightarrow v_{j}$ can be calculated by:
(9) $h_{i}=\begin{cases}c_{v_{i}},&\text{for }v_{i}\in\mathbb{V},\\ W_{h}t_{v_{i}},&\text{otherwise}\end{cases},$
(10) $d=W_{d}v_{j},$
(11) $s_{0}=h_{i}^{T}W_{0}d,\ s_{1}=h_{i}^{T}W_{1}d.$
@@
DocTR (Liao et al., 2023) formulate the entity linking as an anchor word-based entity detection and association problem. Each document entity is represented by the anchor word to convert the entity extraction and linking to token-level tasks. It contains a Deformable DETR-based vision encoder to extract multi-scale visual feature extraction. A LayoutLM based language encoder is applied to encode token-level textual representations. The outputs from vision/language encoders are fed into the vision-language decoder with the language-conditional queries to conduct entity extraction and linking. The decoder queries are one-to-one mapping with language encoder inputs. The entity extraction task aims to predict whether the query underlying token level input is an anchor word and corresponding categories, while entity linking is acquired by
3.2.3. Summary
@@
Several models like Chargrid (Katti et al., 2018) and ACP (Palm et al., 2019) have enhanced VRDU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilised by (Zhang et al., 2020; Wang et al., 2021a), aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialised models, which produce less representative features compared to those generated by large-scale pretrained models like LayoutLM (Xu et al., 2020a) and SelfDoc (Li et al., 2021b). Documents typically exhibit specific layouts and logical structures, which has prompted many models (Yu et al., 2021; Lee et al., 2022b, 2023) to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot (Wang and Shang, 2022) and zero-shot (Wang et al., 2023a) frameworks, which leverage contrastive learning and innovative attention mechanisms. Furthermore, prompt learning has been applied to distil implicit knowledge from large-scale layout-aware pretrained models (Hong et al., 2021; Tu et al., 2023) and large language models (LLMs/MLLMs) (He et al., 2023; Perot et al., 2023). Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRDU optimization.
3.3. VRD Question Answering
@@
Unlike key information extraction, which targets specific details within document images, answering natural language questions involves interpreting more complex intentions and requires models to facilitate interactive understanding between the queries and document representations (Ding et al., 2022). The introduction of DocVQA (Mathew et al., 2021) marked a significant shift in focus from natural scene images to text-dense, layout-aware single-page document images, establishing a benchmark in the field. As advancements have continued, demands have recently emerged for models capable of addressing more complex, multi-page scenarios (Tito et al., 2023; Ding et al., 2023b, 2024a). These emerging requirements highlight the need for models to process multimodal inputs and navigate through extensive documents, reflecting user inquiries’ evolving complexity and naturalness in document-based question-answering systems. This section will briefly review the SoTAs in single-page document VQA models and introduce some recently proposed multi-page document understanding solutions (Tito et al., 2023; Blau et al., 2024).
3.3.1. Single-page VRD-QA
@@
Similar to key information extraction from visually rich documents, single-page question answering (QA) often utilises classical pretrained language models (Devlin, 2018; Liu et al., 2019b) as baselines. These models conduct span-based question answering to extract sequences of text tokens. Additionally, general domain visual language pretrained models (Li et al., 2019; Tan and Bansal, 2019; Kim et al., 2021) are employed to identify the target document’s semantic entities (Ding et al., 2023b). Beyond these plain text or general domain vision language, numerous layout-aware models specifically pretrained in the document domain (see Section 3) have achieved state-of-the-art (SoTA) performance in single-page document tasks. This approach underscores the importance of integrating layout awareness in the preprocessing stages to enhance performance on these specific document-based tasks.
3.3.2. Multi-page VRD-QA
As demand increases for retrieving answers from multi-page documents (Tito et al., 2023; Ding et al., 2024a), most SoTA models (Xu et al., 2020a; Wang et al., 2022b; Hong et al., 2021), which are typically designed for single-page inputs, face a significant challenge due to their input length limitation of 512 tokens. To overcome this limitation, recent innovations have introduced solutions such as transformers capable of handling longer sequences and page-locating modules. These advancements are specifically designed to address the requirements of multipage document question answering (QA), enabling more efficient processing.
Hi-VT5 (Tito et al., 2023) proposes a multimodal hierarchical encoder-decoder architecture for multi-page generative question-answering. A T5-based multimodal transformer is applied to encode the single-page level information, including questions, OCR-extracted page content, image patches, and a set of page tokens. The question and OCR extracted token sequence is the following (Biten et al., 2022) to acquire the layout-aware initial textual representations. Document Image Trasnfoer (DIT) (Li et al., 2022b) is leveraged to acquire initial patch representations. Then, the concatenated question ($Q$) and page OCR tokens ($T$), image patch tokens ($I$), and randomly initialised page tokens ($P$) are fed into T5-based page encoder to enhance the multi-level and multimodal contextual learning contextually. The enhanced page token embeddings $P^{\prime}$ of each input document image are fed into a T5-based decoder to generate the predicted answer, as well auto-regressively, and the page located on the target answer must output the page number based on $P^{\prime}$. As T5 is not a layout-aware language model, masked language modelling is applied to predicted masked tokens by leveraging non-based visual and layout information to boost multimodal understanding. As T5 can accept input sequence lengths up to 20,480 tokens, it dramatically increases the standard VRDU pretrained models that apply Hi-VT5 to multi-page scenarios.
GRAM (Blau et al., 2024): proposes a framework to extend single-page models to tackle multi-page document VQA scenarios. Following the way of pretrained DocFormerv2 (Appalaraju et al., 2024), the framework encodes single-page inputs, including questions, OCR-extracted content, and visual features. For multi-page scenarios, a slim global encoder follows each single-page encoder layer, allowing the new learnable page tokens to interact with other pages contextually. This enables page tokens to capture page-level information through self-attention and enhance document-level understanding with sparse attention. As the global layer is a new stream applied to the pretrained model, an attention bias is applied following ALiBi (Press et al., 2021) to prevent the model from possibly disregarding the page tokens, enabling them to capture more fine-grained information from pretrained weights. During the decoding stage, unlike Hi-VT5, which only feeds page tokens to the decoder, GRAM uses all fine-trained information for the decoder. C-Former (Raffel et al., 2020) is applied to alleviate the high computation consumption, which could revise the information across all pages and distil only important details.
Kang et al. (Kang et al., 2024) introduce a multi-page document VQA framework to use a scoring self-attention mechanism to select and identify the related pages for generating the answer to the input question. The training processes include single-page Document VQA training and then training a self-attention scoring module based on the frozen trained encoder to feed the most relevant page information into the decoder for answer generating. Pix2Struct (Lee et al., 2022a) is used as the single-page model fine-tuned on the DocVQA (Mathew et al., 2021) dataset. The output from the Pix2Struct is fed into the self-attention scoring module to extract the first token for predicting a question-page matching score. The page with the highest matching score will be fed into the decoder for answer generation.
3.3.3. Summary

### v2

@@ 3.1.6 Summary of Key Information Extraction
Several models like Chargrid [chargrid] and ACP [palm2019attend] have enhanced VRD-CU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilized by [trie, ephoie], aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialized models, which produce less representative features compared to those generated by large-scale pre-trained models like LayoutLM [layoutlm] and SelfDoc [selfdoc]. Documents typically exhibit specific layouts and logical structures, which has prompted many models [pick, formnet, formnetv2] to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot [laser] and zero-shot [queryform] frameworks, which leverage contrastive learning and innovative attention mechanisms. Additionally, prompt learning has been applied to distill implicit knowledge from large-scale layout-aware pre-trained models [bros, layoutmask] and large language models (LLMs/MLLMs) [icld3ie, lmdx]. Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRD-CU optimization.
3.2 Document Entity Linking
@@
3.2.1 Entity-Level Entity Linking
@@
DocStruct [docstruct] is the first entity linking framework that predicts hierarchical relationships between document semantic entities using multimodal features. It extracts the [CLS] token from a BERT-like encoder for textual representation ($T_{e}$) and uses an RNN to encode sequential RoI visual features ($V_{e}$) from ResNet-50. The layout feature ($P_{e}$), a linear projection of the bounding box coordinates $[x_{1},y_{1},\dots,x_{4},y_{4}]$, is concatenated with $T_{e}$. The final entity representation is computed as:
$\alpha=\text{Sigmoid}(W[T_{e},P_{e},V_{e}]+b),\ E=[T_{e};P_{e}]+\alpha V_{e},$ (4)
@@
where $\alpha$ gates the influence of visual features, following [wang2019words]. For each entity pair $(E_{i},E_{j})$, the probability of a parent-child link is $P_{i\rightarrow j}=E_{i}ME_{j}$, where $M$ is an asymmetric parameter matrix. Negative sampling [mikolov2013distributed] is used during training to address sparsity and class imbalance.
@@
SERA (Semantic Entity Relation extraction As dependency parsing) [zhang2021entity] frames entity linking as a dependency parsing task. It uses LayoutLM [layoutlm] to extract entity-level textual features $T_{e}$, concatenated with projected label embeddings $L_{e}$ to form entity representations $E=[T_{e},L_{e}]$. Contextual encoders (e.g., Transformer, BiLSTM, Graph) process $E$, and a Biaffine parser computes relation scores:
$h_{i}^{key}=\sigma(W^{key}E_{i}+b^{key}),\ h_{j}^{value}=\sigma(W^{value}E_{j}+b^{value}),\ p_{b}=h_{i}^{key}W_{b1}h_{j}^{value}+h_{i}^{key}W_{b2}.$ (5)
To incorporate layout, a 2D layout feature $l_{ij}$ (min width/height distance) produces a layout score $p_{l}=W_{l}l_{ij}+b_{l}$. The final relation score is $p=p_{b}+p_{l}$, used for binary or multi-label classification.
KVPFormer [kvpformer] reformulates entity linking as a QA task. It uses a Transformer-based encoder-decoder to model joint-grained representations. Each entity’s embedding $E=[T;l]$ is formed by averaging token-level features $T$ and concatenating with its label embedding $l$. A Transformer encoder with spatial attention bias models entity relations. A binary classifier first detects key entities (questions), which are passed to a DETR-style decoder [carion2020end] to predict related entities (answers). For each key entity, top-K answer candidates are ranked by a sigmoid score and refined via a softmax-based coarse-to-fine strategy.
3.2.2 Word-level Entity Linking
@@
As acquiring entity information needs prior knowledge from manual annotation or layout analysing models, some works utilise serialised OCR-extracted sequence of words as inputs to extract the structured relations. However, as the logical relation links semantic entities, word-level frameworks must group word tokens into entities before exploring their association.
@@
Carbonell et al. [carbonell2021named] introduces a framework comprising three modules for sequentially conducting token grouping, entity labelling and relation prediction. Firstly, each text token $t$ is represented by $[L_{t};T_{t}]$ where $L_{t}=[x,y,w,h]$ is the coco format bounding box coordinates of $t$ and $T_{t}$ is the work/representations. All tokens are fed into a token grouping GNN, $\mathcal{G}_{group}$, as a node where the edge between nodes is determined by k-NN to avoid high consumption of fully connected GNN. The $\mathcal{G}_{group}$ is trained on a link prediction task to predict the edge score between two nodes to group words, of which scores larger than a predefined threshold $\rho$. Then, the grouped words are fed into a Graph Attention Network (GAT) to use multi-head attention to aggregate words into entities and follow an MLP to conduct node classification to predict the category of each document entity. At last, another link prediction GNN, $\mathcal{G}_{link}$, is trained on edge classification based on aggregated entities.
@@
SPADE [hwang2021spatial] formulate the word-level entity linking as a spatial dependency parsing task for serializing (ordering and grouping) tokens and predicting inter-group relation between grouped tokens. Firstly, a spatial text encoder is designed to make spatial-aware attention by introducing a relative spatial vector considering relative, physical and angle aspects. During this task, two binary matrices must be predicted $M_{g}$ for token grouping and inter-group linking $M_{l}$. The vertices comprised by a entity types $\mathbb{V}$ and sequence of tokens $\mathbb{T}$ the encoded entity category and token are represented as $c$ and $t$, respectively, and the relation score between vertices $v_{i}\rightarrow v_{j}$ can be calculated by:
$h_{i}=\begin{cases}c_{v_{i}},&\text{for }v_{i}\in\mathbb{V},\\ W_{h}t_{v_{i}},&\text{otherwise}\end{cases},$ (6)
$d=W_{d}v_{j},$ (7)
$s_{0}=h_{i}^{T}W_{0}d,\ s_{1}=h_{i}^{T}W_{1}d.$ (8)
@@
DocTR [doctr] formulate the entity linking as an anchor word-based entity detection and association problem. Each document entity is represented by the anchor word to convert the entity extraction and linking to word-level tasks. It contains a Deformable DETR-based vision encoder to extract multi-scale visual feature extraction. A LayoutLM based language encoder is applied to encode word-level textual representations. The outputs from vision/language encoders are fed into the vision-language decoder with the language-conditional queries to conduct entity extraction and linking. The decoder queries are one-to-one mapping with language encoder inputs. The entity extraction task aims to predict whether the query underlying token level input is an anchor word and corresponding categories, while entity linking is acquired by
3.2.3 Summary of Entity Linking Models
@@
Several models like Chargrid [chargrid] and ACP [palm2019attend] have enhanced VRD-CU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilised by [trie, ephoie], aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialised models, which produce less representative features compared to those generated by large-scale pretrained models like LayoutLM [layoutlm] and SelfDoc [selfdoc]. Documents typically exhibit specific layouts and logical structures, which has prompted many models [pick, formnet, formnetv2] to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot [laser] and zero-shot [queryform] frameworks, which leverage contrastive learning and innovative attention mechanisms. Furthermore, prompt learning has been applied to distil implicit knowledge from large-scale layout-aware pretrained models [bros, layoutmask] and large language models (LLMs/MLLMs) [icld3ie, lmdx]. Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRD-CU optimization.
3.3 VRD Question Answering
@@
Unlike key information extraction, which targets specific details within document images, answering natural language questions involves interpreting more complex intentions and requires models to facilitate interactive understanding between the queries and document representations [vdoc]. The introduction of DocVQA [docvqa] marked a significant shift in focus from natural scene images to text-dense, layout-aware single-page document images, establishing a benchmark in the field. As advancements have continued, demands have recently emerged for models capable of addressing more complex, multi-page scenarios [mpdocvqa, pdfvqa, mmvqa]. These emerging requirements highlight the need for models to process multimodal inputs and navigate through extensive documents, reflecting user inquiries’ evolving complexity and naturalness in document-based question-answering systems. This section will briefly review the SoTAs in single-page document VQA models and introduce some recently proposed multi-page document understanding solutions [mpdocvqa, gram].
3.3.1 Single-page VRD-QA
@@
Similar to key information extraction, single-page question answering (QA) on VRDs often begins with classical pretrained language models such as BERT [bert] and RoBERTa [roberta], which perform span-based QA by extracting relevant text token sequences. In addition, general-domain vision-language models like VisualBERT [visualbert], LXMERT [lxmert], and ViLT [vilt] have been adopted to identify semantic entities in documents [pdfvqa]. Beyond these plain-text or general-domain visual-language models, numerous layout-aware architectures—specifically pretrained on document-centric data (see Section 3)—have achieved state-of-the-art performance on single-page document QA tasks. These approaches highlight the critical role of incorporating layout information during preprocessing to improve model performance on visually structured documents.
3.3.2 Multi-page VRD-QA
With the growing need to retrieve answers from multi-page documents [mpdocvqa, mmvqa], existing single-page models [layoutlm, lilt, bros] face limitations due to their 512-token input constraints. Recent advancements address this by introducing long-sequence transformers, hierarchical encoders, and page-locating modules tailored for multi-page VRD-QA.
Hi-VT5 [mpdocvqa] proposes a hierarchical encoder-decoder framework based on T5 [t5] for multi-page generative QA. Each page is encoded independently using a T5-based multimodal encoder, incorporating layout-aware question and OCR tokens [biten2022latr], visual patch features from DIT [dit], and learnable page tokens. The enhanced page token representations ($P^{\prime}$) are aggregated and passed to the decoder to generate answers. As Hi-VT5 uses T5 (capable of handling up to 20,480 tokens), it significantly expands the input capacity beyond traditional VRD models. To compensate for T5’s lack of inherent layout awareness, masked language modeling is employed during pretraining to inject visual and spatial signals.
GRAM [gram] builds on a pretrained single-page model (DocFormerV2 [docformerv2]) and introduces a lightweight global encoder between layers to enable inter-page interaction. Learnable page tokens attend across pages via sparse self-attention, capturing document-level context. To align with the pretrained backbone, ALiBi [press2021train] attention bias is applied to preserve attention to these new tokens. Unlike Hi-VT5, which only decodes from page tokens, GRAM incorporates all fine-grained token embeddings during decoding. It also employs C-Former [t5] to compress cross-page information, reducing computational overhead.
Kang et al. [kang2024multi] propose a two-stage framework using Pix2Struct [pix2struct] as a single-page encoder fine-tuned on DocVQA [docvqa]. A self-attention-based scoring module is trained on top of the frozen encoder to select the most relevant page based on question-page matching. The selected page is then fed into the decoder to generate the answer, ensuring efficient retrieval by focusing only on the most informative content.
Overall, these multi-page VRD-QA frameworks advance beyond single-page limitations by leveraging hierarchical encoding, inter-page attention, and selective page routing to support more complex, large-scale document understanding tasks.
3.3.3 Summary of VRD-QA Models